OpenStudy (anonymous):

Suppose you choose a marble from a bag containing 3 red marbles, 5 white marbles, and 4 blue marbles. You return the first marble to the bag and then choose again. Find P(red and blue)

OpenStudy (anonymous):

help me bro

OpenStudy (kropot72):

First you need to find the individual probabilities of drawing a red marble and a blue marble. \[P(red)=\frac{number\ of\ red\ marbles}{toal\ number\ of\ marbles}\] Can you do that?

OpenStudy (anonymous):

yeahhh bruh thanks

OpenStudy (kropot72):

What is your result?

OpenStudy (anonymous):

7/12

OpenStudy (kropot72):

Not really. \[P(red)=\frac{number\ of\ red\ marbles}{total\ number\ of\ marbles}=\frac{3}{12}\] Now you need to find the probability of drawing a blue marble which is \[P(blue)=\frac{number\ of\ blue\ marbles}{total\ number\ of\ marbles}\] The probability P(red and blue) is given by \[P(red)\times P(blue)=you\ can\ calculate\]